OpenStudy (anonymous):

There are 20 students on the school's student council. A special homecoming dance committee is to be formed by randomly selecting 6 students from student council. How many possible committees can be formed?

OpenStudy (raden):

it means 20C6

OpenStudy (raden):

use the combination formula : nCr = n!/(n-r)!r!

OpenStudy (raden):

do u know about factotial ?

OpenStudy (raden):

*fACTORIAL

OpenStudy (anonymous):

I was confused because it looked like the answer wouldve been 27,907,200 possible committee. I got that by 20 times 6, then kept lowering the number i multiplied by 6 until i got to 15x6

OpenStudy (raden):

let's check it ... 20C6 = 20!/(20-6)!6! = 20!/14!6! = 20*19*18*17*16*15*14!/(14!*6*5*4*3*2*1) = (20*19*18*17*16*15*)/(6*5*4*3*2*1) = 19 * 17 * 8 * 15 = 38,760 ways hope i didnt mistake